OpenStudy (swissgirl):

Prove: For every positive real number x, there is a positive real number y less than x, with the property that for all positive real numbers z, yz is greater or equal to z

OpenStudy (anonymous):

there is no \(x\) in your conclusion

OpenStudy (swissgirl):

There doesnt have to be

OpenStudy (anonymous):

ooh i see

OpenStudy (swissgirl):

It doesnt have to be true. you can provide a counterexample

OpenStudy (asnaseer):

it looks false to me

OpenStudy (asnaseer):

take the case of x=1

OpenStudy (asnaseer):

then y<1

OpenStudy (asnaseer):

and yz < z

OpenStudy (swissgirl):

ok i see

OpenStudy (anonymous):

i think this is not correct

OpenStudy (anonymous):

\[yz\geq z\] and \(y, z>0\implies y\geq 1\)

OpenStudy (anonymous):

and it is certainly not the case that given ANY \(x\) there is a \(y<x\) where \(y\geq 1\)

OpenStudy (anonymous):

for example \(x=\frac{1}{2}\)

OpenStudy (anonymous):

ok if i was paying attention i would have seen @asnaseer wrote this already

OpenStudy (asnaseer):

:) thats alright @satellite73 - your proof is much neater! :)

OpenStudy (anonymous):

must be miller time

OpenStudy (swissgirl):

hahahah THANKS GUYS :))))))

OpenStudy (asnaseer):

yw :)